RC Meeting and Runplan, Tuesday, August 28

From HallCWiki
Jump to: navigation, search
  • Status
    • More work to tune beam into Hall C yesterday evening - did not converge. Little/no beam to Hall C on OWL shift due to other issues.
      • One hypothesis is that one or more quads that were worked over summer have wrong polarity (note: checks during day shift indicate polarities ok)
    • Controlled Access today while RF recovery ongoing
      • IHA3H07A fixed (encoder cable connected)
      • Quad polarity check
      • Walkthrough training
      • SHMS PMT check
    • Target fan stopped during DAY shift - warmed up target a little, fan restarted
  • Run plan
    • Tuesday SWING, Tuesday OWL and DAY
      • Beam optics tuning
      • Big BPM calibration/checkout test plan
      • After initial setup - SHMS steering plan will be executed (please call RC if not already on site). When MCC asks, ramp up magnets (one at a time) to settings for -8.5 GeV.
      • Beam checkout
      • For the following, see detailed run plan here: [1]
        • Detector and trigger checkout. Call experts before start (after hole-target is centered): Malace, Pooser, Berdnikov, Ambrose for detector checkout, Sawatzky, Malace, Pooser, Huber, Trotta for trigger checkout)
        • Calibration runs